5. (a) Suppose X is a gamma variate with E(x) = 3 and var(X ) = .7 Find the parameters 

α and λ of the gamma distribution. 


1
Expert's answer
2021-03-31T14:56:06-0400

A random variable X which has the gamma distribution with a shape parameter "\alpha" and a rate parameter "\lambda" (i.e. "X\sim \Gamma (\alpha ,\lambda )" ) has the following mean and variance:

"E(X)=\alpha/\lambda=3"

"Var(X)=\alpha/\lambda^2=0.7"

Hence

"\alpha=E(X)^2/Var(X)=3^2/0.7=12.857"

"\lambda=E(X)/Var(X)=3/0.7=4.2857"